# deep neural network STFT short-time Fourier transform

### v1

@@ 1 Introduction
RIR play a crucial role in audio signal processing, as they characterize how sound propagates and decays in a given environment [1]. To obtain room impulse responses that characterize natural reverberation, they can either be measured in actual rooms or simulated using established acoustic models [2]. These natural RIRs reflect the physical properties and geometries of actual spaces, yielding reverberation with consistent time decay, energy distribution, and frequency response characteristics. Accordingly, most RIR datasets and dereverberation benchmarks are based on this natural room acoustics paradigm, aiming to improve technologies such as speech recognition, hearing aids, and audio enhancement in real-world environments [3].
However, there is a discrepancy between natural reverberation and reverb effects used in modern music production. In music production, often artificial RIRs are employed using digital reverb audio effect plugins. Unlike natural reverberation, artificial reverbs are unconstrained by the laws of physics and can exhibit a wide range of exotic behaviors: extremely long or modulated decays, non-monotonic frequency responses, non-linear characteristics, and even deliberately unnatural acoustics. These reverberation effects are part of the creative palette, shaping the aesthetics of contemporary music. As a consequence, dereverberation approaches trained and evaluated exclusively on natural RIR datasets may not generalize well to these artificial reverberation scenarios.
@@ 2 Background
While reverb is a powerful tool for enhancing emotion and depth in music, it can severely degrade the quality and intelligibility of speech and vocals. This is particularly problematic for individuals using hearing aids, for whom reverb can create substantial comprehension difficulties [6, 7]. In music production, it is usually preferable to apply artificial reverb to non-reverberant (dry) vocals, since this allows precise parameter control, which requires the removal of existing natural or artificial reverb. Non-reverberant vocals are also preferred for further music production tasks such as remixing a song, as they provide greater flexibility for creative processing.
Removing reverb from a signal, especially from vocals or speech, is a challenging problem in the field of signal processing. Singing voice dereverberation is conceptually similar to speech dereverberation, and similar methods can be used to address both tasks. However, despite their similarities, speech and singing vocals have distinct characteristics, particularly in terms of pitch, power, and phoneme duration.
Early approaches to vocal and music dereverberation primarily relied on signal processing techniques such as linear prediction [8, 9, 10, 11], or source models designed to capture the harmonic structure of musical signals [12, 13]. These methods typically assumed a structured representation of the source and exploited music’s inherent periodicity and tonal characteristics to mitigate reverberation effects.
Recent research in machine learning has introduced a variety of new approaches, including end-to-end deep neural networks [14]. Although initially proposed for reverb conversion, such models can also be applied to dereverberation. However, these supervised methods tend to learn the characteristics of specific reverbs present in the training data, leading to reduced performance when encountering unseen reverb types, a limitation given the wide range of possible reverberation scenarios in real-world applications. To address this challenge, more recent work has increasingly focused on generative approaches, such as diffusion models, which can be supervised or unsupervised and aim to achieve robust generalization across diverse reverb conditions.
@@ 3 ReverbFX Dataset
We release the ReverbFX dataset [^1] https://sp-uhh.github.io/reverbfx, comprising 1,846 RIRs, with RT60 times ranging from 0.31 to 52.08 seconds. A histogram of RT60 times is shown in Figure 1. All RIRs were generated using four publicly available audio plugins, with further details provided in Section 3.1. To measure each RIR, a Dirac impulse was used as the input signal; this was created from a WAV file in which the first sample is set to 1 and all subsequent samples are set to 0. The reverb effects were applied using DawDreamer [17], a Python API that provides essential DAW functionalities, including integration of Steinberg VST plugins.
For each plugin, we utilized both factory and custom presets. To further increase variation in the resulting audio effects, we generated 14 randomized versions of each preset by modifying two randomly selected parameters from a predefined subset (e.g., delay time and low cut). This procedure results in a total of 15 impulse responses per preset. For all configurations, the dry/wet parameter was explicitly set to 1 to ensure a fully wet signal.
A rigorous validation procedure was employed to filter out anomalous or corrupted data to guarantee the quality and reliability of the RIRs. Each RIR underwent the following checks:
@@
  • RT60 Consistency: Retaining only those RIRs with a computed RT60 in the plausible range of 0.1 to 100 seconds.
@@
  • Early Energy Concentration: Assessing the proportion of energy in the initial segment to ensure it is not overly concentrated at the start.
@@
CSD [20] $4.7$ $1$ f $1810$ Children Songs Korean, English $44.1$ $16$
PJS [21] $0.5$ $1$ m $195$ Children and Anime Songs Japanese $48$ $24$
Opencpop [22] $5.2$ $1$ f $2517$ Chinese Pop Songs Chinese $44.1$ $24$
NUS-48E [23] $1.9$ $6$ m, $6$ f $703$ Traditional and Pop Songs English $44.1$ $16$
NHSS [24] (Test) $4.8$ $5$ m, $5$ f $490$ Pop Songs English $48$ $16$
@@ 3.1 Reverb Audio Effect Plugins
The Protoverb plugin is proprietary software distributed by u-he Heckmann Audio GmbH [^2] https://u-he.com and licensed for free, non-commercial use. Contrary to traditional algorithmic reverbs that aim to minimize resonances, Protoverb builds up as many room resonances as possible, modeling the body of air in a room [25]. It is based on a delay network, whose structure is controlled by a text-based randomization feature, allowing users to generate unique reverb settings. For our study, we created $200$ presets with different random strings in the reverb section to capture a diverse set of RIRs
@@
SkyNet Reverb by Shift Line  [^3] https://shift-line.com is proprietary software offered under a Pay-What-You-Want model.
The SkyNet plugin is based on the two reverb algorithms SkyNet-1 and SkyNet-2 from the Astronaut space reverb pedal, and is designed to create expansive soundscapes with long decay times and diffuse sound [26]. The plugin offers a parameter to control the ratio between the two algorithms.
@@
The Tal-Reverb-4 Plugin is proprietary software distributed by TAL Software GmbH [^4] https://tal-software.com. It is based on a modulated delay network and is designed to evoke the character of 1980s digital reverbs [27]. The algorithm produces a high-quality, plate-style reverb with a diffuse and modulated character, making it particularly suitable for vocals and synthetic sounds.
@@
The Valhalla Supermassive plugin is proprietary software distributed by Valhalla DSP [^5] https://valhalladsp.com and offered as freeware. Designed for extreme and lush reverb and delay effects, it implements a modulated FDN architecture optimized for massive spatialization, long decay times, and evolving textures [28]. The plugin offers a range of algorithm, each with unique temporal and spectral characteristics, making it particularly suited for ambient sound design and experimental processing.
@@ 4.1 Data
Inspired by Lemercier et al. [29], we utilize a variety of singing voice datasets, including OpenSinger [18], M4Singer [19], CSD [20], PJS [21], OpenCpop [22], NUS-48E [23], and NHSS [24] These datasets feature studio-quality recordings with balanced gender and phoneme distributions and include vocals in multiple languages. The combined vocal data amounts to a total of 94 hours. An overview of the datasets and their key characteristics is provided in Table 1.
@@
For every split, each vocal recording is paired with a randomly selected RIR from the corresponding subset. To ensure compatibility for convolution, vocal files are first resampled to 48 kHz. Each RIR is truncated at its maximum index to remove trailing silence or noise, then normalized, converted to mono, and convolved with the vocal signal.
@@ 4.2 Baselines
In a follow-up study [30], the authors adapted the model for 48 kHz audio and conducted an extensive hyperparameter search to optimize the noise schedule and amplitude compression of the input representation. We adopt these optimal hyperparameters to train SGMSE+ on SingingReverbFX using the official implementation. [^6] https://github.com/sp-uhh/sgmse As an ablation, we also train SGMSE+ on natural RIRs, using the datasets employed in [30], to illustrate the domain shift between natural and artificial reverberation.
@@
Jucic et al. [5] extended SGMSE+ by employing a tractable Schrödinger bridge as the diffusion process for speech enhancement and dereverberation. The Schrödinger bridge transforms the input speech distribution into the clean speech distribution by formulating and solving an optimal transport problem. Richter et al. [31, 32] later adapted this model for 48 kHz audio and integrated it into their code repository, which we use for training the Schrödinger bridge on SingingReverbFX. We utilize the stochastic differential equation (SDE) sampler, as it has been shown to provide superior performance for dereverberation tasks [5].
@@ 4.3 Metrics
  • ESTOI [35]: Measures intelligibility by comparing short-time spectral features between reference and output, giving a score from 0 (unintelligible) to 1 (perfect).
@@ 5 Results
In Table 2, we present the singing voice dereverberation results obtained using SingingReverbFX. It can be seen that all methods yield improvements over the reverberant test set across all metrics. Furthermore, training SGMSE+ on natural reverb (with mismatched RIRs) leads to a performance decrease compared to training SGMSE+ with in-domain data from the SingingReverbFX training set. The Schrödinger bridge also demonstrates strong performance, but is slightly outperformed by SGMSE+. This may be attributed to the extensive hyperparameter tuning applied to the 48 kHz model in [30].

### v2

@@ 1 Introduction
RIR play a crucial role in audio signal processing, as they characterize how sound propagates and decays in a given environment [1]. To obtain room impulse responses that characterize natural reverberation, they can either be measured in actual rooms or simulated using established acoustic models [2]. These RIRs reflect the physical properties and geometries of actual spaces, yielding reverberation with consistent time decay, energy distribution, and frequency response characteristics. Accordingly, most RIR datasets and dereverberation benchmarks are based on this natural room acoustics paradigm, aiming to improve technologies such as speech recognition, hearing aids, and audio enhancement in real-world environments [3].
However, there is a discrepancy between natural reverberation and artificial reverb effects used in modern music production [4, 5]. Unlike real-world acoustics, music production often relies on digital reverb effects that are directly applied to individual tracks, such as vocals, using software plugins. These plugins can employ convolution-based methods using recorded or synthesized RIRs, or algorithmic approaches that simulate reverberation through mathematical models. Artificial reverbs are unconstrained by the laws of physics and can exhibit a wide range of exotic behaviors: extremely long or modulated decays, non-monotonic frequency responses, non-linear characteristics, and even deliberately unnatural acoustics. These reverberation effects are part of the creative palette, shaping the aesthetics of contemporary music. As a consequence, dereverberation approaches trained and evaluated exclusively on natural RIR datasets may not generalize well to these artificial reverberation scenarios.
@@ 2 Background
Music datasets often provide audio in separate stems, such as vocals, drums, bass, and accompaniment. These stems can be used for training source separation models [8], stem mixing and mastering [9], and other music information retrieval tasks [10]. However, vocal stems are often provided with added reverb, and truly anechoic (dry) voice recordings are not always available. Therefore, singing voice dereverberation is necessary to retrieve the original anechoic voice. In turn, the resulting dry singing voice can be convolved with a variety of RIRs to generate additional data for training purposes–a common data augmentation strategy. Moreover, in music production, it is usually preferable to apply reverb to non-reverberant vocals, as this allows precise parameter control and greater creative flexibility.
Removing reverb from a signal, especially from vocals or speech, is a challenging problem in the field of signal processing. Singing voice dereverberation is conceptually similar to speech dereverberation, and similar methods can be used to address both tasks. However, despite their similarities, speech and singing voice have distinct characteristics, particularly in terms of pitch, power, and phoneme duration [11].
Early approaches to vocal and music dereverberation primarily rely on signal processing techniques such as linear prediction [12, 13, 14, 15], or source models designed to capture the harmonic structure of musical signals [16, 17]. These methods typically assume a structured representation of the source and exploit music’s inherent periodicity and tonal characteristics to mitigate reverberation effects.
Recent research in machine learning has introduced a variety of new approaches, including end-to-end deep neural networks [18]. These supervised methods tend to learn the characteristics of specific reverbs present in the training data. This leads to reduced performance when encountering unseen reverb types, a limitation given the wide range of possible reverberation scenarios in real-world applications. To address this challenge, more recent work has increasingly focused on generative approaches, such as diffusion models, which can be supervised or unsupervised and aim to achieve robust generalization across diverse reverb conditions [19].
@@ 3 ReverbFX Dataset
We release the ReverbFX dataset [^1] https://sp-uhh.github.io/reverbfx, comprising 1,846 RIRs, with RT60 times ranging from 0.31 to 52.08 seconds. A histogram of RT60 times is shown in Figure 1. All RIRs are generated using four publicly available audio plugins (see Section 3.1) and measured using a discrete-time unit impulse as the input signal. The reverb effects are applied via DawDreamer [22], a Python API providing DAW functionality, including support for VST plugins.
For each plugin, we utilize both factory and custom presets. To further increase variation in the resulting audio effects, we generate 14 randomized versions of each preset by modifying two randomly selected parameters from a predefined subset (e.g., delay time and low cut). This procedure results in a total of 15 impulse responses per preset. For all configurations, the dry/wet parameter is set to 1 to ensure a fully wet signal.
A rigorous validation procedure is employed to filter out anomalous or corrupted data to guarantee the quality and reliability of the RIRs. Each RIR underwent the following checks:
@@
  • RT60 Consistency: Retaining only those RIRs with a computed RT60 in the range of 0.1 to 100 seconds.
@@
CSD [25] $4.7$ 11 f $1810$ Children Songs Korean, English $44.1$ $16$
PJS [26] $0.5$ 11 m $195$ Children and Anime Songs Japanese $48$ $24$
Opencpop [27] $5.2$ 11 f $2517$ Chinese Pop Songs Chinese $44.1$ $24$
NUS-48E [28] $1.9$ 66 m, $6$ f $703$ Traditional and Pop Songs English $44.1$ $16$
NHSS [29] (Test) $4.8$ 55 m, $5$ f $490$ Pop Songs English $48$ $16$
@@ 3.1 Reverb Audio Effect Plugins
The Protoverb plugin is proprietary software distributed by u-he Heckmann Audio GmbH [^2] https://u-he.com and licensed for free, non-commercial use. Contrary to traditional algorithmic reverbs that aim to minimize resonances, Protoverb builds up as many room resonances as possible, modeling the body of air in a room [30]. It is based on a delay network, whose structure is controlled by a text-based randomization feature, allowing users to generate unique reverb settings. For our study, we created $200$ presets with different random strings in the reverb section to capture a diverse set of RIRs.
@@
SkyNet Reverb by Shift Line [^3] https://shift-line.com is proprietary software offered under a Pay-What-You-Want model. It is based on the two reverb algorithms SkyNet-1 and SkyNet-2 from the Astronaut space reverb pedal, and is designed to create expansive soundscapes with long decay times and diffuse sound [31]. The plugin offers a parameter to control the mix ratio between the two algorithms.
@@
Tal-Reverb-4 is a proprietary software distributed by TAL Software GmbH [^4] https://tal-software.com. It is based on a modulated delay network and is designed to evoke the character of 1980s digital reverbs [32]. The algorithm produces a high-quality, plate-style reverb with a diffuse and modulated character, making it particularly suitable for vocals and synthetic sounds.
@@
The Valhalla Supermassive plugin is proprietary software distributed by Valhalla DSP [^5] https://valhalladsp.com and is offered as freeware. Designed for extreme and lush reverb and delay effects, it implements a modulated FDN architecture optimized for massive spatialization, long decay times, and evolving textures [33]. The plugin offers a range of algorithms, each with unique temporal and spectral characteristics, making it particularly suited for ambient sound design and experimental processing.
@@ 4.1 Data
Inspired by Lemercier et al. [34], we utilize a variety of singing voice datasets, including OpenSinger [23], M4Singer [24], CSD [25], PJS [26], OpenCpop [27], NUS-48E [28], and NHSS [29]. These datasets feature studio-quality recordings with balanced gender and phoneme distributions and include vocals in multiple languages. The combined vocal data amounts to a total of 94 hours. An overview of the datasets and their key characteristics is provided in Table 1.
@@
For every split, each vocal recording is paired with a randomly selected RIR from the corresponding subset. To ensure compatibility for convolution, vocal files are first resampled to 48 kHz. Each RIR is truncated at its maximum index to remove trailing silence or noise, then normalized, converted to mono by taking the left channel, and convolved with the vocal signal.
@@ 4.2 Baselines
In a follow-up study [35], the authors adapted the model for 48 kHz audio and conducted an extensive hyperparameter search to optimize the noise schedule and amplitude compression of the input representation. We adopt these optimal hyperparameters to train SGMSE+ on SingingReverbFX using the official implementation. [^6] https://github.com/sp-uhh/sgmse As an ablation, we also train SGMSE+ on natural RIRs using the datasets employed in [35] to illustrate the domain shift between natural and artificial reverberation.
@@
Jukić et al. [7] extended SGMSE+ by employing a tractable Schrödinger bridge as the diffusion process for speech enhancement and dereverberation. The Schrödinger bridge transforms the input speech distribution into the clean speech distribution by formulating and solving an optimal transport problem. Richter et al. [36, 37] later adapted this model for 48 kHz audio and integrated it into their code repository, which we use for training the Schrödinger bridge on SingingReverbFX. We utilize the stochastic differential equation (SDE) sampler, as it has been shown to provide superior performance for dereverberation tasks [7].
@@ 4.3 Metrics
  • ESTOI [40]: Measures intelligibility by comparing short-time spectral features between the reference and output, giving a score from 0 (unintelligible) to 1 (perfect).
@@ 5 Results
In Table 2, we present the singing voice dereverberation results obtained using SingingReverbFX. It can be seen that all methods yield improvements over the reverberant test set across all metrics. Furthermore, training SGMSE+ on natural reverb (with mismatched RIRs) leads to a performance decrease compared to training SGMSE+ with in-domain data from the SingingReverbFX training set. The Schrödinger bridge also demonstrates strong performance but is slightly outperformed by SGMSE+. This may be attributed to the extensive hyperparameter tuning applied to the 48 kHz model in [35].